2. Alleles D and L are codominant alleles that define fur color in Madagascar lemurs. Lemurs that are homozygous for allele D (genotype DD) have dark brown fur. Lemurs that are homozygous for the L allele (genotype LL) have very light brown fur. The DL heterozygotes are generally light brown, but they have a dark patch or a collar around their neck. Among the 600 Lemurs that you studied on Madagascar, the distribution of genotypes was as follows: Dark brown (DD)-40 individuals Light brown with dark patches (DL)-160 individuals Light brown (LL)-400 individuals. Find the frequency of each allele in the population fr(L) = 3. Cystic fibrosis is a recessive genetic disorder caused by allele c. People with genotypes CC and Cc for this trait are not affected by the disease. In modern Germany, the frequency of allele c is 0.02 a) What is the frequency of dominant allele C? fr(C) Assume that the population does not deviate from Hardy-Weinberg equilibrium. Estimate the frequencies of the following genotypes: fr(Cc)= fr(CC)- 4. The recessive allele 1 codes for lactose intolerance. People who are recessive homozygotes (11) cannot digest milk; dominant homozygotes (LL) and heterozygotes (LI) have no problem digesting milk. Among 400 people from Eastern Mongolia, 36 were unable to digest milk. hat this population does not deviate from the Hardy-Weinberg equilibrium, estimate a. Assuming th the frequency of the L and 1 alleles: fr(1) fr(L) = b. Find the frequency of each of the following genotypes fr(lI)- fr(LI)-
Answer #1
  • Dominance can be complete or incomplete or codominance.

Answer:

2)

  • In complete dominance one gene expression is completely suppressed by the other.
  • In incomplete dominance neither the dominant nor the recessive but an intermediate character between the two characters is expressed.
  • In codominance both the characters are expressed in an interspersed manner.

Given that,

lemurs with DD = 40

lemurs with DL = 160

lemurs with LL = 400

total number of lemurs = 600

frequency of allele D = proportion of lemurs with DD + 1/2 proportion of lemurs with DL

= 40/600 + 1/2 (160/600)

= 0.066 + 0.133

= 0.199 = 0.2 approximately

frequency of L allele = proportion of lemurs with LL + half the proportion of lemurs with DL

= 400/600 + 1/2 (160/600)

= 0.666 + 0.133

= 0.799 = 0.8 approximately

Therefore the frequency of allele D in above population is 0.2 and that of allele L is 0.8.

2) given data,

C is dominant over c

frequency of c = 0.02

and the population is in Hardy Weinberg equilibrium

from Hardy Weinberg principle,

p (frequency of allele C) + q (frequency of allele c) = 1

p + 0.02 = 1

p = 1 - 0.02

p = 0.98

again from Hardy Weinberg principle

p2(frequency of genotype CC) + 2pq (frequency of genotype Cc) + q2 (frequency of genotype cc) = 1

from this,

frequency of genotype CC, p2 = 0.98 x 0.98 = 0.96

frequency of genotype Cc , 2pq = 2 x 0.98 x 0.02 = 0.0392

frequency of genotype cc, q2 = 0.02 x 0.02 = 0.0004

Therefore frequency of allele C is 0.98, frequencies of genotypes CC, Cc and cc are 0.96, 0.0392 and 0.0004 respectively if the frequency of allele c is 0.02.